# Local geometry and quantum geometric tensor of mixed states

### v4

@@ Abstract
The quantum geometric tensor (QGT) is a fundamental concept for characterizing the local geometry of quantum states. After casting the geometry of pure quantum states and extracting the QGT, we generalize the geometry to mixed quantum states via the density matrix and its purification. The gauge-invariant QGT of mixed states is derived, whose real and imaginary parts are the Bures metric and the Uhlmann form, respectively. In contrast to the imaginary part of the pure-state QGT that is proportional to the Berry curvature, the Uhlmann form vanishes identically for ordinary physical processes. Moreover, there exists a Pythagorean-like equation that links different local distances and reflect the underlying fibration. The Bures metric of mixed states is shown to reduce to the corresponding Fubini-Study metric of the ground state as temperature approaches zero, establishing a correspondence despite the different underlying fibrations. We also present two examples with contrasting local geometries and discuss experimental implications.
@@ III.4 The Uhlmann metric
It can be verified that under the gauge transformation $W^{\prime}=W\mathcal{U}$, $\gamma_{\mu\nu}$, $\sigma_{\mu\nu}$, and $\omega$ change according to
Using this, the transformation (51) becomes
A candidate of a gauge-invariant modification to $\gamma_{\mu\nu}$ is
@@
which is referred to as the Uhlmann metric. For the imaginary part $\sigma_{\mu\nu}$, the transformation (52) can be reformulated into
The expression motivates a gauge-invariant modification
@@
Using Eq. (53), it can be verified that both $g^{\text{U}}_{\mu\nu}$ and $\sigma^{\text{U}}_{\mu\nu}$ are invariant under the gauge transformation $W^{\prime}=W\mathcal{U}$. The proof is outlined in Appendix F. Their relations to the pure-state counterparts will be studied later.
@@ III.5 Uhlmann distance
The Bures metric can be written as
$g^{\text{B}}_{\mu\nu}=\frac{1}{2}\sum_{ij}\frac{(\sqrt{\lambda_{i}}+\sqrt{\lambda_{j}})^{2}}{\lambda_{i}+\lambda_{j}}\langle i|\partial_{\mu}\sqrt{\rho}|j\rangle\langle j|\partial_{\nu}\sqrt{\rho}|i\rangle$ (63)
according to Eq. (149). Similarly, by expanding the commutator, the Uhlmann connection is expressed as
@@
A calculation of Eq. (III.5) shows that $g^{\text{U}}_{\mu\nu}$ reduces to (see Appendix G for details)
$=g^{\text{B}}_{\mu\nu}$ (67)
according to Eq. (63).
Consequently, the Uhlmann distance is indeed equal to the Bures distance. Eq. (III.5) thus takes the form
@@
With the help of purified states, the definition $\ddot{\text{o}}$qvist distance can also be cast into the form similar to Eq. (43). The purifications of $\rho(t)$ is written as $W(t)=\sum_{n}\sqrt{\lambda_{n}(t)}|n(t)\rangle\langle n(t)|U(t)$. The key to the derivation is to choose the phase factor as $U(t)=\sum_{n}\mathrm{e}^{\mathrm{i}\theta_{n}(t)}|n(t)\rangle\langle n|$, which is unitary due to the facts $U(t)U^{\dagger}(t)=\sum_{n}|n(t)\rangle\langle n(t)|=1$ and $U^{\dagger}(t)U(t)=\sum_{n}|n\rangle\langle n|=1$. Thus, the purification is $W(t)=\sum_{n}\sqrt{\lambda_{n}(t)}|n(t)\rangle\langle n|\mathrm{e}^{\mathrm{i}\theta_{n}(t)}$, whose corresponding purified state is $|W(t)\rangle=\sum_{n}\sqrt{\lambda_{n}(t)}|n(t)\rangle\otimes\mathrm{e}^{\mathrm{i}\theta_{n}(t)}|n\rangle$. Using the expressions, Eq. (69) can be written as
@@ III.7 Uhlmann form
The proof is summarized in Appendix G. We emphasize that the proof and result can be generalized situations where the restriction of $\sigma^{\text{U}}$ on $\mathcal{D}^{N}_{N}$ can be relaxed. This is done by replacing $A_{\text{U}}$ by $\pi^{*}A_{\text{U}}$, a one-form on the total space $\mathcal{S}_{N}$, in the expression above to get
@@ IV Correspondence between pure and mixed states
The fibrations $S^{2N-1}/\text{U}(1)=CP^{N-1}$ and $\mathcal{S}_{N}/\text{U}(N)=\mathcal{D}^{N}_{N}$ lay the foundation for the fiber-bundle descriptions of the Berry and Uhlmann phases, respectively. While the bundle corresponding to the Berry phase may be nontrivial, the Uhlmann bundle is always trivial [61, 50, 44, 50]. Consequently, the Uhlmann connection and Uhlmann curvature do not reduce to the Berry connection and Berry curvature as $T\rightarrow 0$ because they belong to fiber bundles with distinct topologies. Nevertheless, it has been conditionally proven that the Uhlmann phase approaches the Berry phase as $T\rightarrow 0$, which is also known as the Uhlmann-Berry correspondence [45]. One may ask if there exists any other correspondence between the geometric results of pure and mixed states? Here we show that for systems of dimension $N>1$, the Bures metric of mixed states indeed reduces to the Fubini-Study metric of pure states as $T\rightarrow 0$, thereby providing a correspondence. To our knowledge, there has been no proof of the correspondence between the two metrics. Ref. [64] provides an example based on qubits but requires a constant factor to be dropped. Here we provide a general proof showing that the correspondence is exact.
@@
The normalized eigenstate $|i\rangle$ satisfies $\langle i|\mathrm{d}i\rangle+\langle\mathrm{d}i|i\rangle=0$, resulting in $\langle i|\mathrm{d}\sqrt{\rho}|i\rangle=\mathrm{d}\sqrt{\lambda_{i}}$. A straightforward evaluation shows (see Appendix H for details)
@@ V.1 Generic result for two-level systems
Figure 1: Bures metric of the spin-$\frac{1}{2}$ paramagnet as a function of temperature: (Top panel) $g^{\text{B}}_{\theta\theta}$, (Bottom panel) $g^{\text{B}}_{\phi\phi}$ at $\theta=\frac{\pi}{4}$.
@@ V.2 Spin-$\frac{1}{2}$ system
Figure 2: Bures metric of the toy model as a function of temperature, (Up) $g^{\text{B}}_{xx}$, (Bottom) $g^{\text{B}}_{xy}$, where $k_{x}=k_{y}=1.2\pi$ and $\mu=1.0$.
@@ VII Acknowledgments
We thank Prof. Guang-Yu Guo for bringing our attention to the QGT. H.G. was supported by the National Natural Science Foundation of China (Grant No. 12074064). C.C.C. was supported by the National Science Foundation under Grant No. PHY-2310656 and thanks the NCTS, Taiwan for its hospitality during his visit. X. Y. H. was supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611).
@@ Appendix C Properties of the space of mixed states
We have denoted by $\mathcal{P}$ the space of all density matrices. For $\rho_{1}$, $\rho_{2}\in\mathcal{P}$ and a pair of arbitrary complex numbers $\lambda$ and $\mu$, it is possible to find counterexamples to $\lambda\rho_{1}+\mu\rho_{2}\in\mathcal{P}$. On the other hand, as a complex matrix, $\rho$ is also equivalent to an operator acting on the $N$-tuple vectors in $\mathcal{H}=\mathbb{C}^{N}$. Consequently, $\rho$ belongs to $End(\mathcal{H})$, which is the space of operators on $\mathcal{H}$, and is also known as the algebra of $N\times N$ complex matrices. This means $\mathcal{P}\subset End(\mathcal{H})$. Furthermore, the requirement that $\text{Tr}(\lambda\rho_{1}+\mu\rho_{2})=1$ leads to $\lambda+\mu=1$, which further implies $\mathcal{P}$ is a convex subset of $End(\mathcal{H})$ [2].
@@ Appendix F Modification of the real and imaginary parts of $\gamma_{\mu\nu}$
Under the gauge transformation $W^{\prime}=W\mathcal{U}$, the second term on the right-hand-side of Eq. (56) changes as
@@
where we have applied $\partial_{\mu,\nu}\mathcal{U}\mathcal{U}^{\dagger}=-\mathcal{U}\partial_{\mu,\nu}\mathcal{U}^{\dagger}$, $\mathcal{U}^{\dagger}\partial_{\nu}\mathcal{U}\mathcal{U}^{\dagger}=-\partial_{\nu}\mathcal{U}^{\dagger}$ and the cyclic property of the trace. Using Eqs. (III.4) and (F), it is straightforward to verify that $g^{\text{U}\prime}_{\mu\nu}=g^{\text{U}}_{\mu\nu}$.
@@
Appendix G Details of the Uhlmann metric and Uhlmann form
@@
Appendix H Proof of Eq. (IV)

### v5

@@ Abstract
The quantum geometric tensor (QGT) is a fundamental concept for characterizing the local geometry of quantum states. After casting the geometry of pure quantum states and extracting the QGT, we generalize the geometry to mixed quantum states via the density matrix and its purification. The gauge-invariant QGT of mixed states is derived, whose real and imaginary parts are the Bures metric and the Uhlmann form, respectively. In contrast to the imaginary part of the pure-state QGT that is proportional to the Berry curvature, the Uhlmann form vanishes identically for ordinary physical processes. Moreover, there exists a Pythagorean-like equation that links different local distances and reflect the underlying fibration. The Bures metric of mixed states is shown to reduce to the corresponding Fubini-Study metric of the ground states as temperature approaches zero, establishing a correspondence despite the different underlying fibrations. We also present two examples with contrasting local geometries and discuss experimental implications.
@@ III.4 The Uhlmann metric
It can be verified that under the gauge transformation $W^{\prime}=W\mathcal{U}$, $\gamma_{\mu\nu}$, $\sigma_{\mu\nu}$, and $\omega$ change accordingly. A candidate of a gauge-invariant modification to $\gamma_{\mu\nu}$ is
@@
which is referred to as the Uhlmann metric. Moreover, a gauge-invariant modification of the imaginary part leads to
@@
One can verify that both $g^{\text{U}}_{\mu\nu}$ and $\sigma^{\text{U}}_{\mu\nu}$ are invariant under the gauge transformation $W^{\prime}=W\mathcal{U}$. The proof is outlined in Appendix F. Their relations to the pure-state counterparts will be studied later.
@@ III.5 Uhlmann distance
As summarized in Appendix E, the Bures metric can be written as
$g^{\text{B}}_{\mu\nu}=\frac{1}{2}\sum_{ij}\frac{(\sqrt{\lambda_{i}}+\sqrt{\lambda_{j}})^{2}}{\lambda_{i}+\lambda_{j}}\langle i|\partial_{\mu}\sqrt{\rho}|j\rangle\langle j|\partial_{\nu}\sqrt{\rho}|i\rangle.$ (57)
Similarly, by expanding the commutator, the Uhlmann connection is expressed as
@@
We show in Appendix G that the Uhlmann distance is indeed equal to the Bures distance. Eq. (III.5) thus takes the form
@@ III.7 Uhlmann form
The proof is summarized in Appendix G. We emphasize that the proof and result can be generalized to situations where the restriction of $\sigma^{\text{U}}$ on $\mathcal{D}^{N}_{N}$ can be relaxed. This is done by replacing $A_{\text{U}}$ by $\pi^{*}A_{\text{U}}$, a one-form on the total space $\mathcal{S}_{N}$, in the expression above to get
@@
Moreover, the vanishing of the Uhlmann form also highlights the different topologies of pure and mixed states. Here we refer to the U(1) principal bundle of pure states that admits a Berry connection as the Berry bundle, which can be topologically nontrivial. Quantized topological invariants, such as the Chern number, can be found by calculating the characteristic classes of the bundle. In contrast, the Uhlmann bundle is topologically trivial [61, 50, 44, 50], and its characteristic classes must vanish. Since the Uhlmann form is gauge-invariant, its integral over $\mathcal{D}^{N}_{N}$ may serve as a topological invariant of the Uhlmann bundle, similar to the Chern number of the Berry bundle. If the integral is non-zero, it would contradict the topological triviality of the Uhlmann bundle. We also remark that the Uhlmann bundle cannot reduce to the Berry bundle as $T\rightarrow 0$ because the former requires full-rank density matrices. Since topological properties are defined with respect to the corresponding geometry, the difference between the $T\rightarrow 0$ and $T=0$ cases comes from the change of the underlying bundles. However, there still exists some correspondence between quantities derived from pure and mixed states as $T\rightarrow 0$. which will be discussed in the next section. Nevertheless, our findings show that $\mathcal{D}^{N}_{N}$ already possesses rich geometric properties for mixed states. We summarize our main results by comparing the key points between pure and mixed states in Table 1.
@@ IV Correspondence between pure and mixed states
The fibrations $S^{2N-1}/\text{U}(1)=CP^{N-1}$ and $\mathcal{S}_{N}/\text{U}(N)=\mathcal{D}^{N}_{N}$ lay the foundation for the fiber-bundle descriptions of the Berry and Uhlmann phases, respectively. As mentioned previously, the Berry bundle can be topologically nontrivial, whereas the Uhlmann bundle is always trivial. Consequently, the Uhlmann connection and Uhlmann curvature do not reduce to the Berry connection and Berry curvature as $T\rightarrow 0$ because they belong to fiber bundles with distinct topologies. Nevertheless, it has been conditionally proven that the Uhlmann phase approaches the Berry phase as $T\rightarrow 0$, which is also known as the Uhlmann-Berry correspondence [45]. One may ask if there exists any other correspondence between the geometric results of pure and mixed states? Here we show that for systems of dimension $N>1$, the Bures metric of mixed states indeed reduces to the Fubini-Study metric of pure states as $T\rightarrow 0$, thereby providing a correspondence. To our knowledge, there has been no proof of the correspondence between the two metrics. Ref. [64] provides an example based on qubits but requires a constant factor to be dropped. Here we provide a general proof showing that the correspondence is exact.
@@
The normalized eigenstate $|i\rangle$ satisfies $\langle i|\mathrm{d}i\rangle+\langle\mathrm{d}i|i\rangle=0$, resulting in $\langle i|\mathrm{d}\sqrt{\rho}|i\rangle=\mathrm{d}\sqrt{\lambda_{i}}$. A straightforward evaluation shows (see Appendix I for details)
@@ V.1 Generic result for two-level systems
Figure 1: Bures metric of the spin-$\frac{1}{2}$ paramagnet as a function of temperature: (Top panel) $g^{\text{B}}_{\theta\theta}$, (Bottom panel) $g^{\text{B}}_{\phi\phi}$ at $\theta=\frac{\pi}{4}$. The dots at $T=0$ show the corresponding values from the Fubini-Study metric.
@@ V.2 Spin-$\frac{1}{2}$ system
Figure 2: Bures metric of the toy model as a function of temperature, (Up) $g^{\text{B}}_{xx}$, (Bottom) $g^{\text{B}}_{xy}$, where $k_{x}=k_{y}=1.2\pi$ and $\mu=1.0$. The dots at $T=0$ show the corresponding values from the Fubini-Study metric.
@@ VII Acknowledgments
We thank Prof. Guang-Yu Guo for bringing our attention to the QGT. H.G. was supported by the National Natural Science Foundation of China (Grant No. 12074064) and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904). C.C.C. was supported by the National Science Foundation under Grant No. PHY-2310656 and thanks the NCTS, Taiwan for its hospitality during his visit. X. Y. H. was supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611).
@@ Appendix C Properties of the space of mixed states
We have denoted by $\mathcal{P}$ the space of all density matrices. For $\rho_{1}$, $\rho_{2}\in\mathcal{P}$ and a pair of arbitrary complex numbers $\lambda$ and $\mu$, it is possible to find counterexamples to $\lambda\rho_{1}+\mu\rho_{2}\in\mathcal{P}$. On the other hand, as a complex matrix, $\rho$ is also equivalent to an operator acting on the $N$-tuple vectors in $\mathcal{H}=\mathbb{C}^{N}$. Consequently, $\rho$ belongs to $End(\mathcal{H})$, which is the space of operators on $\mathcal{H}$, and is also known as the algebra of $N\times N$ complex matrices. This means $\mathcal{P}\subset End(\mathcal{H})$. Furthermore, the requirement that $\text{Tr}(\lambda\rho_{1}+\mu\rho_{2})=1$ leads to $\lambda+\mu=1$, which further implies $\mathcal{P}$ is a convex subset of $\mathcal{P}\subset End(\mathcal{H})$ [2].
@@ Appendix F Modification of the real and imaginary parts of $\gamma_{\mu\nu}$
Under the gauge transformation $W^{\prime}=W\mathcal{U}$,
Using this, the transformation (139) becomes
Moreover, the second term on the right-hand-side of Eq. (51) changes as
@@
where we have applied $\partial_{\mu,\nu}\mathcal{U}\mathcal{U}^{\dagger}=-\mathcal{U}\partial_{\mu,\nu}\mathcal{U}^{\dagger}$, $\mathcal{U}^{\dagger}\partial_{\nu}\mathcal{U}\mathcal{U}^{\dagger}=-\partial_{\nu}\mathcal{U}^{\dagger}$ and the cyclic property of the trace. Using Eqs. (F) and (F), it is straightforward to verify that $g^{\text{U}\prime}_{\mu\nu}=g^{\text{U}}_{\mu\nu}$.
For the imaginary part $\sigma_{\mu\nu}$, the transformation (140) can be reformulated into
@@
Appendix G Details of Uhlmann metric, Bures metric, and Uhlmann form
A calculation of Eq. (III.5) shows that $g^{\text{U}}_{\mu\nu}$ reduces to (see the proof below)
$=g^{\text{B}}_{\mu\nu}$ (149)
according to Eq. (57).
@@ Appendix G Details of Uhlmann metric, Bures metric, and Uhlmann form
Appendix H Details $\ddot{\text{o}}$qvist distance
[…]
With the help of purified states, the definition $\ddot{\text{o}}$qvist distance can also be cast into the form similar to Eq. (43). The purifications of $\rho(t)$ is written as $W(t)=\sum_{n}\sqrt{\lambda_{n}(t)}|n(t)\rangle\langle n(t)|U(t)$. The key to the derivation is to choose the phase factor as $U(t)=\sum_{n}\mathrm{e}^{\mathrm{i}\theta_{n}(t)}|n(t)\rangle\langle n|$, which is unitary due to the facts $U(t)U^{\dagger}(t)=\sum_{n}|n(t)\rangle\langle n(t)|=1$ and $U^{\dagger}(t)U(t)=\sum_{n}|n\rangle\langle n|=1$. Thus, the purification is $W(t)=\sum_{n}\sqrt{\lambda_{n}(t)}|n(t)\rangle\langle n|\mathrm{e}^{\mathrm{i}\theta_{n}(t)}$, whose corresponding purified state is $|W(t)\rangle=\sum_{n}\sqrt{\lambda_{n}(t)}|n(t)\rangle\otimes\mathrm{e}^{\mathrm{i}\theta_{n}(t)}|n\rangle$.
Appendix I Proof of Eq. (IV)